Circular No. 6819

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE P. Garnavich, Harvard-Smithsonian Center for Astrophysics (CfA), re-

ports that the High-Z Supernova Search Team (IAUC 6160, 6646) has discovered nine supernovae on CCD images taken with the Cerro Tololo Interamerican Observatory (CTIO) 4-m telescope by R. Schommer (CTIO), B. Schmidt (Mount Stromlo and Siding Springs Observatory), and S. Jha and P. Challis (CfA). The supernovae were identified by subtracting images taken on 1997 Dec. 29–30 UT from those obtained on 1998 Jan. 23–24. The candidates were confirmed with spectra and images taken on Jan. 25–26 and Feb. 1 with the Keck-2 telescope by A. V. Filippenko, A. G. Riess, and D. C. Leonard (University of California, Berkeley).

SN  1998 UT  α_{2000}  δ_{2000}  I  z type
1998F Jan. 23 4^{h}16^{m}50.^{s}13 − 5^{o}44′59′′.6 24.5 0.52 ?
1998G Jan. 23 8 03 37.02 + 6 10 13.9 22.8 0.30 II?
1998H Jan. 23 8 04 51.47 + 5 36 39.3 23 0.66 ?
1998I Jan. 23 8 04 51.56 + 5 15 47.7 23.6 0.89 Ia
1998J Jan. 23 9 31 10.48 − 4 45 36.5 22.5 0.83 Ia
1998K Jan. 24 4 13 42.86 − 5 50 45.2 23.8 ?  ?
1998L Jan. 24 11 33 36.63 + 4 35 04.6 23.6 ?  II?
1998M Jan. 24 11 33 44.37 + 4 05 13.4 23.3 0.63 Ia
1998N Jan. 24 11 33 29.39 + 3 51 12.5 23.1 0.26 ?

Each supernova is within 2′′ of its host galaxy’s center. A foreground galaxy with z = 0.02 is centered 2′′.5 southeast of SN 1998M. Finder charts may be requested by sending e-mail to pgarnavich@cfa.harvard.edu.

4U 1608–52 W. Cui, Center for Space Research, Massachusetts Institute of Tech-

nology; and J. Swank, Goddard Space Flight Center, report on behalf of the ASM team at MIT and RXTE Science Operation Facility: “The real-time ASM lightcurve indicates that 4U 1608–52, a recurrent soft x-ray transient source, started an x-ray outburst on Jan. 29. Rough daily-averaged 1.3–12keV fluxes: Jan. 29, 25 mCrab; 31, 21; Feb. 1, 28; 2, 121; 3, 261; 4, 491. A pointed observation was carried out with the PCA and HEXTE detectors aboard RXTE on Feb. 3.82 UT, and the measured flux was consistent with the ASM results. More pointed RXTE observations have been planned. Observations at other wavelengths are encouraged.”

1998 February 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green